OpenStudy (anonymous):

A box contains 7 green, 5 yellow & 6 purple balls. Find the probability of yellow or purple ball in a single draw.

OpenStudy (anonymous):

im guessing for yellow it will be: 5/18 ?

OpenStudy (anonymous):

You have a total of 18 balls. So the probability of drawing only the yellow ball is: 5/18. And the probability of drawing only the purple ball it's: 6/18=1/3.

OpenStudy (anonymous):

can you help me with one more problem? :)

OpenStudy (anonymous):

But if it's an "or" statement, then make sure to add them together so you answer the question. And sure =)

OpenStudy (anonymous):

so what would be the final answer? 6/18?

OpenStudy (anonymous):

No 6/18 is only the probability of finding the purple balls. Make sure you add yellow and purple to get the total probability of drawing a yellow OR a purple ball.

OpenStudy (anonymous):

next question: Probability of an event is 10/11, what are the odds that the event will occur?

OpenStudy (anonymous):

so 11/18 ? :)

OpenStudy (anonymous):

yyeeeep. And garr odds. Odds are just the flipped fraction of finding the probability. So if the probability of an event, for example, it 4/6, the odds will be 6/4. Do the same with the number 10/11. =D

OpenStudy (anonymous):

11/10 ? :)

OpenStudy (anonymous):

yush. :)

OpenStudy (anonymous):

thanks! and sorry, but can you check one problem that i did, and im not sure if it's right

OpenStudy (anonymous):

lol np, and sure!

OpenStudy (anonymous):

Nine people are entered in a race. If there are no ties, in how many ways can the first three places come out? I got 472

OpenStudy (anonymous):

Hmm. How did you get that?

OpenStudy (anonymous):

idk, i multiplied random #'s..

OpenStudy (anonymous):

I honestly don't know how to solve this o_O Since there's 3 places, there should be 3 people in that. So that means that 27 people could go into those three places....27 * 9 is 243 I believe...idk lol i don't wanna give you the wrong answer but that's what i think.

OpenStudy (anonymous):

ohh okay!